I need help with this problem. Find an equation of the line containing the given pair of points. Express your answer in the form x=a, y=b, or y=mx+b
(0,0) and (4,7)
what is an equation of the line?
y=__________
Thank you in advance for your help

Answer by stanbon(75887)

(0,0) and (4,7)
what is an equation of the line?
-------------------
slope = (7-0)/(4-0) = 7/4
-------------------
EQUATION form is y = mx+b
Substitute x=0, y=0 and m=(7/4) to solve for b
-------------------
0 = (7/4)*0 + b
b = 0
---------------
EQUATION:
y = (7/4)x
===============
Cheers,
Stan H.